# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 4.2.1 SEMs with missing data
This section evaluates the accuracy of the HVB method for estimating the SEMs with missing values in the response variable using simulated data. For each simulated dataset, missing values in the response variable y are introduced according to the MNAR mechanism using a logistic regression model defined in Section 2.3. To construct this model, an additional covariate $\mathbf{x}^{*}$ is generated from a log-normal distribution with with a log-scale mean of 0 and a log-scale standard deviation of 1, and the model is regressed on $\mathbf{x}^{*}$ and the response variable y. We set $\psi_{0}=1$ (intercept), $\psi_{\textbf{x}^{*}}=-1$ (the coefficient for $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient for y). This configuration results in approximately $40\%$ missing values, leading to $n_{u}=1,752$.
We fit SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models for the simulated datasets with missing values. As $n$ and $n_{u}$ are large, the HVB-AllB algorithm (Algorithm S2) described in Section S4.2 of the online supplement is used. The tuning parameters for the HVB-AllB algorithm is set as follows: the number of MCMC iterations ($N_{1}$) is set to 10, and the block size ($k^{*}$) is defined as $10\%$ of $n_{u}$, resulting in 10 or 11 blocks.
The initial values for the variational parameters corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$ are set as described in Section 4.1, while $\nu$ is initialised with a value of 4. The initial values for the missing data ($\mathbf{y}_{u}$) are simulated from the conditional distribution $p(\mathbf{y}_{u}\mid\boldsymbol{\xi}^{(0)},\mathbf{y}_{o})$, where $\boldsymbol{\xi}^{(0)}$ represents the vector of initial parameter values $\boldsymbol{\xi}$.
@@
We ran the HVB-AllB algorithm for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau models across both simulated datasets. For the SEM-t and YJ-SEM-t models, more iterations were required to achieve convergence due to the additional complexity introduced by estimating the latent variables $\boldsymbol{\tau}$ and the degrees of freedom parameter $\nu$. Specifically, for simulated dataset 1, SEM-t was run for 20,000 iterations and YJ-SEM-t for 10,000 iterations. For simulated dataset 2, SEM-t required 30,000 iterations, while YJ-SEM-t was run for 20,000 iterations.
Table 4 presents the posterior means and 95% credible intervals for some of the model parameters of the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models obtained using the HVB-AllB method applied to the simulated dataset 1 with missing values, along with the computational cost for one VB iteration. The table also includes the $\text{DIC}_{5}$ values for each model, calculated using the formula in Equation (20). Corresponding posterior density plots of the parameters are given in Section S7.2.2 of the online supplement.
We start by comparing the estimated values of the fixed effects (specifically $\beta_{0}$ and ${\beta}_{1}$ for the SEM and $\psi_{0}$ and ${\psi}_{1}$ for the missing data model), the variance parameter $\sigma_{\textbf{e}}^{2}$, and the spatial autocorrelation parameter $\rho$, as these parameters are common to all four models. The posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, $\psi_{1}$, and $\rho$ are nearly identical for both YJ-SEM-Gau and YJ-SEM-t, closely matching the true values. In contrast, the estimates from SEM-Gau and SEM-t show significant deviations from the true values. The estimated posterior means of $\sigma_{\textbf{e}}^{2}$ from SEM-Gau, SEM-t, and YJ-SEM-Gau differ considerably from the true value, while the posterior mean from YJ-SEM-t is much closer to the true value. The models incorporating the YJ transformation (YJ-SEM-Gau and YJ-SEM-t) successfully recover the parameter $\gamma$. The estimates of $\nu$ obtained from the SEM-t and YJ-SEM-t models are slightly inaccurate. Furthermore, the posterior means of the estimated missing values from YJ-SEM-t and YJ-SEM-Gau align more closely with the true missing values compared to those from SEM-Gau and SEM-t; see Figure S7 in Section S7.2.2 of the online supplement.
For the simulated dataset 1 with missing values in the response variable, the YJ-SEM-t model provides the best fit, as indicated by the lowest $\text{DIC}_{5}$ value, among all four models. This result is expected, as the dataset was generated from the YJ-SEM-t model, which incorporates both heavy tails and skewness. The second-best performer is YJ-SEM-Gau, which has the next lowest DIC value. Furthermore, similar to the case without missing values discussed in Section S6 of the online supplement, the relatively small difference in the $\text{DIC}_{5}$ values between YJ-SEM-t and YJ-SEM-Gau suggests that YJ-SEM-Gau achieves a model fit nearly comparable to that of YJ-SEM-t for this dataset.
The left panel of Figure 4 shows the kernel densities of the true missing values ($\mathbf{y}_{u}$) and the kernel density of posterior means of the missing values at locations $s_{1},s_{2},\ldots,s_{n}$, estimated from different SEMs for the simulated dataset 1. The densities from both YJ-SEM-Gau and YJ-SEM-t are nearly identical and closely align with the density of the true missing values.
The right panel of Figure 4 presents the posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) predicted by each SEM. The true value of the maximum missing value lies closer to the posterior densities produced by YJ-SEM-Gau and YJ-SEM-t than to those from SEM-Gau and SEM-t. These outcomes are as expected, since YJ-SEM-Gau and YJ-SEM-t yield similar parameter estimates and DIC values (see Table 4). Figure S8 in Section S7.2.2 of the online supplement provides similar density plots for the simulated dataset 2.
Figure 4: Left panel: The kernel density of the true missing values ($\textbf{y}_{u}$) and the kernel density of posterior means of the missing values ($\textbf{y}_{u}$) obtained using the HVB-AllB method for different SEMs for the simulated dataset 1 with missing values. Right panel: The posterior density of the maximum missing value ($\textbf{y}_{u_{\textrm{max}}}$) obtained by different SEMs for the simulated dataset 1 with missing values. The true value of the maximum missing value is indicated by the vertical line.
@@
Table 4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 with missing data, along with the computation time (CT) in seconds for one HVB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Table 5 reports the posterior means and $95\%$ credible intervals for key parameters of the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, obtained using the HVB-AllB method on the simulated dataset 2 with missing values, along with the computational cost for one VB iteration. The table also presents the corresponding $\text{DIC}_{5}$ values for each model. Posterior density plots for these parameters are provided in Section S7.2.2 of the online supplement. The estimated posterior means for $\beta_{0}$, $\beta_{1}$, $\psi_{0}$, and $\sigma^{2}_{\textbf{e}}$ across all four models closely match their true values. Similarly, the estimated posterior mean of $\gamma$ for the YJ-SEM-Gau and YJ-SEM-t models is also close to the true value. The parameter $\nu$ is accurately estimated in both the SEM-t and YJ-SEM-t models. In addition, all four models produce posterior means for the estimated missing values that closely match the true values; see Figure S8 in Section S7.2.2 of the online supplement.
@@
Table 5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 with missing data, along with the computation time (CT) in seconds for one HVB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
Based on the $\text{DIC}_{5}$ values, the YJ-SEM-t model offers the best fit for the simulated dataset 2 with missing values. In contrast, for the simulated dataset 2 without missing values, the SEM-Gau model provides the best fit; see Section S6 of the online supplement. For the simulated dataset 2 with missing data, the SEM-t model has the second-lowest $\text{DIC}_{5}$, indicating it is the second-best performing model.
As shown in Tables 4 and 5, the HVB-AllB algorithms for SEM-t and YJ-SEM-t, which incorporate Student’s $t$ errors, require significantly more computation time than those for SEM-Gau and YJ-SEM-Gau. This increased computational cost arises from the greater complexity of VB (HVB-AllB) optimisation in models with $t$ errors. Specifically, these models introduce an additional length $n$ latent variable vector and include the degrees of freedom parameter $\nu$, both of which increase the dimensionality and complexity of the VB optimisation.
Our simulations demonstrate that the standard Gaussian VB method for full data and the hybrid VB (HVB-NoB and HVB-AllB) methods for missing data offer accurate yet computationally efficient approximations for estimating both Gaussian and non-Gaussian SEMs, outperforming HMC in terms of computational cost. The YJ-SEM-Gau and YJ-SEM-t models are particularly effective for modeling data with pronounced skewness and heavy tails. In most scenarios, the DIC values show that YJ-SEM-Gau performs nearly as well as YJ-SEM-t. Although estimating YJ-SEM-t incurs significantly higher computational cost compared to SEM-Gau (the standard SEM with Gaussian errors), YJ-SEM-Gau has a similar computational cost to SEM-Gau while offering substantial improvements in modeling non-Gaussian data.
@@ 5 Real data application
We apply the proposed SEMs and VB methods to analyse a dataset of 25,357 single-family homes sold between 1993 and 1998, along with a spatial weight matrix of size $357\times 25,$ 357 × 25 , 357 from Lucas County, Ohio, USA. This dataset is available in the R package spData (Bivand et al.,, 2023). The data set includes various house characteristics: house age, the lot size in square feet, the number of rooms, the total living area in square feet, the number of bedrooms, and binary indicators for each year from 1993 to 1998 to represent the year of house sale. For our analysis, we focus on the house price data from 1998, which contains $4,378$ observations. We call this dataset Lucas-1998-HP. House prices are typically positive and often exhibit skewness. To transform house prices into the entire real line and reduce skewness, researchers commonly apply logarithmic transformations before modeling (LeSage and Pace,, 2004; Wijayawardhana et al.,, 2024). Accordingly, we apply the natural logarithmic transformation to the house prices in the Lucas-1998-HP dataset. Despite this transformation, the resulting variable still exhibits left skewness (see Figure S9 in Section S8 of the online supplement).
To evaluate the impact of missing data on Gaussian and non-Gaussian SEMs, we use the Lucas-1998-HP dataset, a complete real-world dataset. We first fit the SEMs to the full dataset using the VB algorithm. Then, we intentionally introduce missing values into the response variable and refit the SEMs using the proposed HVB-AllB method. This setup allows for a direct comparison of parameter estimates obtained with and without missing data.
For the Lucas-1998-HP dataset, we fit all four models described in Section 2. The set of independent variables includes various powers of house age (age, $\text{age}^{2}$, and $\text{age}^{3}$), the natural logarithm of the lot size in square feet ($\ln(\text{lotsize})$), the number of rooms (rooms), the natural logarithm of the total living area in square feet ($\ln(\text{LTA})$), and the number of bedrooms. The natural logarithm of house price in hundreds of thousands is selected as the dependent variable. Results for the complete dataset (without missing values) are presented in Section 5.1, while Section 5.2 summarises the findings with missing data.
@@ 5.1 Full Data SEMs
This section presents the results of analysing the Lucas-1998-HP dataset by fitting all four SEMs introduced in Section 2, using the VB algorithms outlined in Section S2 of the online supplement. The starting values for the VB algorithms are obtained as follows: The variational mean vector ($\boldsymbol{\mu}$) is set to the estimates for $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE). The initial value for $\nu$ is set to 4, and the initial value for $\gamma$ is set to 1. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors.
As in the simulation study with full data described in Section S6 of the online supplement, models with Student-$t$ errors required more iterations to achieve convergence. The VB algorithms were run for 20,000 iterations for the SEM-Gau and YJ-SEM-Gau models and 75,000 iterations for the SEM-t and YJ-SEM-t models, at which point convergence was attained. Convergence was assessed visually by examining the plots of variational means over iterations, which are provided in Section S9.2 of the online supplement.
The estimated posterior means and $95\%$ credible intervals for selected model parameters across different models are summarised in Table 6, along with the corresponding DIC values. According to the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, the YJ-SEM-t model is the most appropriate SEM for the Lucas-1998-HP dataset, followed by the YJ-SEM-Gau model as the second most suitable.
@@
Table 6: Posterior means and 95% credible intervals for selected parameters of different SEMs for the Lucas-1998-HP dataset (no missing values). The table also includes DIC values for each model. Parameters labeled as ’NA’ indicate that they are not applicable to the corresponding model.
@@
We now compare estimates of key parameters across different SEMs. A crucial aspect of SEM estimation is identifying the underlying spatial correlation, represented by the parameter $\rho$. For the SEM-Gau, YJ-SEM-Gau, and YJ-SEM-t models, the estimated values of $\rho$ are relatively close, with posterior mean values of 0.6291, 0.6019, and 0.5990, respectively. This suggests a moderate spatial correlation in the Lucas-1998-HP dataset. However, for the SEM-t model, the estimated $\rho$ is noticeably lower, with a posterior mean value of 0.5475, compared to the other three models.
@@
Similar to linear regression models, the fixed effects $\boldsymbol{\beta}$ in SEMs represent the influence of covariates on the response variable, which in this case is the logarithm of the house price. An interesting disparity arises in the estimated posterior mean of the fixed effect corresponding to the covariate ’age’ ($\beta_{\text{age}}$). For the SEM-Gau model, $\beta_{\text{age}}$ is positive, while for the other three models, $\beta_{\text{age}}$ is negative. This suggests that for the SEM-Gau model, the age of the house is associated with an increase in house prices, while for the other models, the age of the house is linked to a decrease in house prices. The estimated posterior means of the fixed effect for the number of rooms ($\beta_{\text{rooms}}$) are 0.0068, 0.0188, 0.0311, and 0.0238 for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models, respectively. These estimates suggest that the effect of the number of rooms on house prices varies across models. Notably, the YJ-SEM-Gau model shows the strongest positive influence, while the SEM-Gau model exhibits the weakest.
@@
Additionally, the SEM-t, YJ-SEM-Gau, and YJ-SEM-t models provide further insights into the characteristics of the dataset. The posterior mean of $\nu=3.0249$ for SEM-t and $\nu=7.9482$ for YJ-SEM-t suggest that the log-transformed house price distribution exhibits heavy tails. Furthermore, the posterior means of $\gamma$, $1.5264$ for YJ-SEM-Gau and $1.5529$ for YJ-SEM-t, suggest left skewness in the response variable, effectively capturing the asymmetry in the distribution of log-transformed house prices. These distributional features are visually evident in the kernel density plot of the response variable (see Figure S9 in Section S8 of the online supplement).
5.2 SEMs with missing data

### v2

@@ 7.1 SEMs with missing responses
This section evaluates the robustness and accuracy of the proposed SEMs and the HVB method in the presence of missing responses, using simulated datasets 1 and 2. First, missing values are introduced in the response variable y of each simulated dataset under the MNAR mechanism using the logistic regression model in Equation (11). To construct this model, an additional covariate $\mathbf{x}^{*}$ is generated from a log-normal distribution with a log-scale mean of 0 and a log-scale standard deviation of 1, and the model is regressed on $\mathbf{x}^{*}$ and the response variable y. For the dataset 1, the logistic regression parameters are set to $\psi_{0}=1$ (intercept), $\psi_{\textbf{x}^{*}}=-1$ (the coefficient for $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient for y). This configuration results in approximately $40\%$ missing responses, leading to $n_{u}=1,729$. For the dataset 2, the intercept is set to $\psi_{0}=0.5$ (with the other coefficients unchanged), leading to approximately 40% missing responses ($n_{u}=1,774$).
We fit SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to these simulated datasets with missing values. As $n$ and $n_{u}$ are large, the HVB-AllB algorithm, which employs the MCMC steps detailed in Algorithm S2 in Section S5.4 of the online supplement, is utilised. The tuning parameters for the HVB-AllB algorithm are set as follows: the number of MCMC iterations ($N_{1}$) is set to 10, and the block size ($k^{*}$) is defined as $10\%$ of $n_{u}$, resulting in 11 blocks. The initial values for the variational mean vector $\boldsymbol{\mu}$ corresponding to $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, $\rho$, $\gamma$, as well as the initial values for the missing responses $\mathbf{y}_{u}$, are specified using the procedure described in Section 6.2. For models with Student’s $t$ errors, the degrees of freedom parameter $\nu$ and the set of latent variables $\boldsymbol{\tau}=(\tau_{1},\dots,\tau_{n})^{\top}$ must also be initialised. The variational mean corresponding to $\nu$ is initialised at 4. The variational means corresponding to the latent variables are initialised by sampling from the prior distribution $p(\tau_{i}\mid\nu^{(0)})\sim\text{IG}(\frac{\nu^{(0)}}{2},\frac{\nu^{(0)}}{2})$ for $i=1,\dots,n$, where $\nu^{(0)}=4$ denotes the initial value of $\nu$. For the variational covariance matrix, all elements of B and the diagonal elements of D are initialised to 0.01.
@@
We run the HVB-AllB algorithms for 10,000 iterations for the SEM-Gau and YJ-SEM-Gau across both simulated datasets. For the SEM-t and YJ-SEM-t, convergence generally requires more iterations due to the additional complexity introduced by the latent variables $\boldsymbol{\tau}$ and the parameter $\nu$. However, the number of required iterations varies across datasets and models. For simulated dataset 1, SEM-t is run for 20,000 iterations, whereas YJ-SEM-t achieves convergence within 10,000 iterations. For simulated dataset 2, SEM-t requires 30,000 iterations, while YJ-SEM-t is run for 20,000 iterations. Convergence diagnostic plots are provided in Section S11.2 of the online supplement.
7.1.1 Results for the simulated dataset 1
[…]
Table 4: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 1 with missing data, along with the computation time (CT) in seconds for one VB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
@@ 7.1.1 Results for the simulated dataset 1
[…]
The left four panels of Figure 4 compare the kernel density of the true missing values ($\textbf{y}_{u}$) with the kernel densities of the posterior means of the missing values estimated under different SEMs for the simulated dataset 1. The densities from both YJ-SEM-Gau and YJ-SEM-t are nearly identical and closely align with the density of the true missing values. The right panel of Figure 4 presents the posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, estimated under each SEM. The true value of the maximum missing value lies closer to the posterior densities produced by YJ-SEM-Gau and YJ-SEM-t than to those produced by SEM-Gau and SEM-t. This behaviour is expected, as YJ-SEM-Gau and YJ-SEM-t provide parameter estimates that are closer to the true values and yield more comparable $\text{DIC}_{5}$ values (see Table 4).
Figure 4: Left panel: The kernel density of the true missing values ($\textbf{y}_{u}$) and the kernel densities of posterior means of the missing values obtained using the HVB-AllB method for different SEMs for the simulated dataset 1 with missing values. Right panel: The posterior density of the maximum missing value, $\textrm{max}(\textbf{y}_{u})$, obtained by different SEMs for the simulated dataset 1 with missing values. The true value of the maximum missing value is indicated by the vertical line.
7.1.2 Results for the simulated dataset 2
[…]
Table 5: Posterior means and 95% credible intervals for selected parameters of different SEMs based on the simulated dataset 2 with missing data, along with the computation time (CT) in seconds for one VB iteration. The table also presents the $\text{DIC}_{5}$ values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
@@ 7.1.2 Results for the simulated dataset 2
Based on the $\text{DIC}_{5}$ values, the YJ-SEM-t provides the best fit for simulated dataset 2 with missing values, while the SEM-t has the second-lowest $\text{DIC}_{5}$, indicating it is the second-best performing model.
[…]
7.2 Summary and computation time comparison
This simulation study evaluates how well the proposed SEMs capture key characteristics of non-Gaussian spatial data, both with fully observed data and in the presence of missing values, using two simulated datasets: dataset 1 exhibits skewness and heavy tails, while dataset 2 is closer to Gaussian. This section summarises the findings of the simulation study with missing data, while Section S8.3 of the online supplement summarises the corresponding results for the full data.
As shown in Tables 4 and 5, the HVB-AllB algorithm applied to the SEM-t and YJ-SEM-t, both of which assume Student’s $t$-distributed errors, requires more computation time per iteration compared to SEM-Gau and YJ-SEM-Gau. This is due to the added complexity of VB optimisation: these models introduce an additional latent variable vector of length $n$ and include the degrees of freedom parameter $\nu$, increasing the dimensionality of the optimisation. In addition, SEM-t and YJ-SEM-t typically require more VB iterations to converge. Convergence diagnostics plots in Section S11.2 of the online supplement illustrate this behaviour. Consequently, HVB-AllB for SEM-t and YJ-SEM-t is generally more time-consuming than for SEM-Gau and YJ-SEM-Gau, both because each iteration is slower and more iterations are needed to achieve convergence.
In terms of performance, YJ-SEM-Gau and YJ-SEM-t are particularly effective for modelling skewed and heavy-tailed data. For simulated dataset 1, YJ-SEM-t attains the lowest $\textrm{DIC}_{5}$ value, indicating the best fit, while YJ-SEM-Gau achieves the second-lowest value, which is close to that of YJ-SEM-t. For simulated dataset 2, which is closer to Gaussian, all four models recover the true parameter values accurately and yield nearly identical $\textrm{DIC}_{5}$ values, with YJ-SEM-t again achieving the lowest. Although estimating YJ-SEM-t incurs substantially higher computational cost compared to SEM-Gau (the conventional SEM with Gaussian errors), YJ-SEM-Gau maintains a computation time comparable to SEM-Gau while providing considerable improvements when modelling non-Gaussian data. Similar patterns are observed in the analysis of the fully observed datasets, as presented in Section S8 of the online supplement.
@@ 8 Real data application
We apply the proposed SEMs and VB methods to a dataset of single-family homes sold in Lucas County, Ohio, USA. This dataset is available in the R package spData (Bivand et al.,, 2023). The data set includes various house characteristics: house age, the lot size in square feet, the number of rooms, the total living area in square feet, the number of bedrooms, and binary indicators for each year from 1993 to 1998 to represent the year of house sale. For our analysis, we focus on the house price data from 1998, which contains $4,378$ observations. We call this dataset Lucas-1998-HP. House prices are typically positive and often exhibit skewness. To transform house prices into the entire real line and reduce skewness, researchers commonly apply logarithmic transformations before modelling (LeSage and Pace,, 2004; Wijayawardhana et al.,, 2024). Accordingly, we apply the natural logarithmic transformation to the house prices in the Lucas-1998-HP dataset. However, even after the transformation, the resulting log-prices remain left-skewed; see Figure S9 in Section S10 of the online supplement.
First, the proposed VB algorithm (Algorithm S1 in Section S2 of the online supplement) is used to estimate all four SEMs using the full Lucas-1998-HP dataset. Next, missing values are introduced into the response variable, and the SEMs are re-estimated using the proposed HVB-AllB algorithm (Algorithm 1, together with the MCMC steps in Algorithm S2 in Section S5.4 of the online supplement). This design enables a direct comparison between estimates obtained from the full dataset and those obtained in the presence of missing values, with the full-data estimates treated as the ground truth. The set of predictor variables includes various powers of house age (age, $\text{age}^{2}$, and $\text{age}^{3}$), the natural logarithm of the lot size in square feet ($\ln(\text{lotsize})$), the number of rooms (rooms), the natural logarithm of the total living area in square feet ($\ln(\text{LTA})$), and the number of bedrooms (beds). The natural logarithm of house price in hundreds of thousands of US dollars is selected as the response variable. Results for the complete dataset (without missing values) are presented in Section 8.1, while Section 8.2 provides the findings with missing data.
@@ 8.1 Full data SEMs
This section presents the results of fitting all four SEMs to the full Lucas-1998-HP dataset. The initial values for the VB algorithms are set following the same approach used in the simulation studies in Sections 6 and 7, and we use $p=4$ factors. Consistent with the simulation results with full data (see Section S8 of the online supplement), the models with Student’s $t$ errors require more iterations to achieve convergence. Accordingly, the VB algorithms are run for 20,000 iterations for SEM-Gau and YJ-SEM-Gau, and 75,000 iterations for SEM-t and YJ-SEM-t. Convergence is assessed visually using the trajectories of the variational means, with the diagnostic plots provided in Figure S20 in Section S11.3 of the online supplement.
The estimated posterior means and $95\%$ credible intervals for selected model parameters across different SEMs, along with their corresponding DIC values, are summarised in Table 6. The posterior density plots for these parameters are provided in Figure S10 in Section S10 of the online supplement. According to the $\text{DIC}_{1}$ and $\text{DIC}_{2}$ values, the YJ-SEM-t is the most appropriate SEM for the Lucas-1998-HP dataset, followed by the YJ-SEM-Gau as the second most suitable.
Table 6: Posterior means and 95% credible intervals for selected model parameters of different SEMs for the Lucas-1998-HP dataset (without missing values). The table also includes DIC values for each model. Parameters labelled as ’NA’ indicate that they are not applicable to the corresponding model.
@@
We now compare estimates of key parameters across different SEMs. A crucial aspect of SEM estimation is identifying the underlying spatial correlation, represented by the parameter $\rho$. For the SEM-Gau, YJ-SEM-Gau, and YJ-SEM-t, the estimated values of $\rho$ are relatively close, with posterior mean values of 0.6291, 0.6019, and 0.5990, respectively. This suggests a moderate spatial correlation in the Lucas-1998-HP dataset. However, for the SEM-t, the estimated $\rho$ is noticeably lower, with a posterior mean value of 0.5475, compared to the other three models.
@@
Similar to linear regression models, the fixed effects $\boldsymbol{\beta}$ in SEMs represent the influence of covariates on the response variable, which in this case is the logarithm of the house price. An interesting disparity arises in the estimated posterior mean of the fixed effect for the covariate ’age’ ($\beta_{\text{age}}$). For the SEM-Gau, $\beta_{\text{age}}$ is positive, while for the other three models, $\beta_{\text{age}}$ is negative. This discrepancy reflects a key limitation of SEM-Gau: it is misspecified for this dataset because it cannot adequately capture important features of the response variable, such as skewness and heavy tails. As a result, SEM-Gau produces misleading estimates for the effect of house age, suggesting an increase in price rather than the decrease indicated by the more flexible models. Similarly, the estimated posterior means of the fixed effect for the number of rooms ($\beta_{\text{rooms}}$) are 0.0068, 0.0188, 0.0311, and 0.0238 for SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t, respectively, showing that the effect of the number of rooms on house prices varies across models. Notably, YJ-SEM-Gau shows the strongest positive influence, while SEM-Gau exhibits the weakest, further reflecting the consequences of misspecification.
@@
Additionally, the SEM-t, YJ-SEM-Gau, and YJ-SEM-t provide further insights into the characteristics of the dataset. The posterior means of $\nu=3.0249$ for SEM-t and $7.9482$ for YJ-SEM-t suggest that the log-transformed house price distribution exhibits heavy tails. Furthermore, the posterior means of $\gamma$, $1.5264$ for YJ-SEM-Gau and $1.5529$ for YJ-SEM-t, indicate left skewness in the response variable, effectively capturing its asymmetry. These distributional features are visually evident in the kernel density plot of the response variable (see Figure S9 in Section S10 of the online supplement).
8.2 SEMs with missing responses